OpenStudy (hunt123):

If f(x)=25-x^2 and g(x)=x+5 , what is(f/g) ? Write your answer in simplest form.

OpenStudy (unklerhaukus):

(f/g)(x) = f(x) / g(x)

OpenStudy (hunt123):

that would be x^/25, x/5 what do i do with these

OpenStudy (hunt123):

x^2/25

OpenStudy (hunt123):

would it be 4

OpenStudy (unklerhaukus):

f(x) / g(x) = (25-x^2) / x+5

OpenStudy (hunt123):

5-x

OpenStudy (hunt123):

is should be 5-x yes or no

OpenStudy (unklerhaukus):

the numerator : 25-x^2 is a difference of squares A difference of square can be expressed as a^2-x^2 = (a-x)(a+x) because (a-x)(a+x) = a(a+x)-x(a+x) = a^2+ax-ax+x^2 = a^2-x^2

OpenStudy (unklerhaukus):

yes

OpenStudy (hunt123):

now is that in the simplest term?

OpenStudy (hunt123):

or do i have to do -x+5

OpenStudy (unklerhaukus):

yeah the denominator has divided away

OpenStudy (hunt123):

the end answer would be 5-x

OpenStudy (unklerhaukus):

5-x, or -x+5 i prefer the first arrangement, some people might like the other one better the main point is that we got rid of that demoninator

OpenStudy (hunt123):

okay i might have a few more ill need help on

OpenStudy (hunt123):

(f + g)(2) (f − g)(4) (f ÷ g)(2) (f × g)(1)

OpenStudy (unklerhaukus):

(f + g)(x) = f(x) + g(x)

OpenStudy (unklerhaukus):

you can add the functions and then evaluate with x=2, or evaluate the functions with x=2 and then add.

OpenStudy (hunt123):

Pairs -4 7/3 10 0

OpenStudy (hunt123):

can you show me with one and ill try to get the rest

OpenStudy (unklerhaukus):

it might be easiest to find (f + g)(2) as f(2) + g(2)

OpenStudy (hunt123):

that would be 4

OpenStudy (unklerhaukus):

f(2) = 25 - (2)^2

OpenStudy (hunt123):

?

OpenStudy (unklerhaukus):

are the f and g the same as for the original question?

OpenStudy (hunt123):

If f(x) = 2x + 3 and g(x) = x2 − 1, find the values of combining these functions. Match each combined function to its corresponding value.

OpenStudy (unklerhaukus):

oh right

OpenStudy (hunt123):

thats what everything needs to plug into

OpenStudy (unklerhaukus):

so you have (f+g)(2) = f(2)+g(2) f(x) = 2x + 3 g(x) = x^2 − 1 f(2) = 2(2) + 3 g(2) = (2)^2 − 1

OpenStudy (hunt123):

im so lost right now how will we plug the stuff in tho

OpenStudy (hunt123):

which one of these (f + g)(2) (f − g)(4) (f ÷ g)(2) (f × g)(1) will pair up with these -4 7/3 10 0

OpenStudy (unklerhaukus):

we pluged in x=2 into f(x) = 2x + 3 we got f(2) = 2(2) + 3 = 2 × 2 + 3 =

OpenStudy (hunt123):

7

OpenStudy (unklerhaukus):

yeha, now we plug x=2 into g(x) = x^2 - 1 we got g(2) = 2^2 - 1 =

OpenStudy (hunt123):

5

OpenStudy (unklerhaukus):

almost

OpenStudy (hunt123):

7 as well no it would be 6

OpenStudy (unklerhaukus):

g(2) = 2^2 - 1 = 4 - 1 =

OpenStudy (hunt123):

3

OpenStudy (unklerhaukus):

so f(2) = 7, and g(2) = 3 (f+g)(2) = f(2)+g(2) = 7+3 =

OpenStudy (hunt123):

10

OpenStudy (unklerhaukus):

ok. next! (f − g)(4)

OpenStudy (hunt123):

hold on which one = 10?

OpenStudy (unklerhaukus):

yes 10 is 10

OpenStudy (hunt123):

(f + g)(2) (f − g)(4) (f ÷ g)(2) (f × g)(1) which one of thos = 10

OpenStudy (unklerhaukus):

We plug x=4 into f(x) = 2x + 3 we get f(4) = 2(4) + 3 = 2 × 4 + 3 =

OpenStudy (unklerhaukus):

oh the one we just did was (f + g)(2), right?

OpenStudy (hunt123):

okay on to the next one

OpenStudy (hunt123):

11

OpenStudy (hunt123):

is the answer to what you asked

OpenStudy (unklerhaukus):

and what is g(4)?

OpenStudy (hunt123):

gonna have to go through the steps im sorry

OpenStudy (unklerhaukus):

plug x=4, into g(x)

OpenStudy (hunt123):

7

OpenStudy (hunt123):

its not a guess promise

OpenStudy (hunt123):

cause if g(x)=x^-1 it would be 4*2-1

OpenStudy (hunt123):

4*2=8-1=7

OpenStudy (hunt123):

?

OpenStudy (unklerhaukus):

g(x) = x^2 - 1 we get g(4) = 4^2 - 1 = 4 * 4 - 1

OpenStudy (hunt123):

15

OpenStudy (unklerhaukus):

so (f-g)(4) = ?

OpenStudy (hunt123):

-4

OpenStudy (unklerhaukus):

correct

OpenStudy (hunt123):

can we do the next two as well please

OpenStudy (hunt123):

first it will be 7

OpenStudy (unklerhaukus):

f(x) = 2x + 3 g(x) = x^2 − 1 (f ÷ g)(2) f(2) =7

OpenStudy (hunt123):

next will be 5

OpenStudy (hunt123):

so 7/5 right

OpenStudy (unklerhaukus):

wait g(x) = x^2 − 1 g(2) = (2)^2 - 1 = 2 * 2 - 1

OpenStudy (hunt123):

3

OpenStudy (hunt123):

the answer is 7/3

OpenStudy (unklerhaukus):

yes,

OpenStudy (hunt123):

okay next

OpenStudy (hunt123):

5 for the next one

OpenStudy (unklerhaukus):

Now lastly (f × g)(1)

OpenStudy (hunt123):

next is 0

OpenStudy (hunt123):

5*0 is 0 right

OpenStudy (unklerhaukus):

f(x) = 2x + 3 f(1) = 2(1) + 3 = 5 g(x) = x^2 − 1 g(1) = 1^1 -1 = 0 (f × g)(1) = 5 * 0 = 0 \(\checkmark\)

OpenStudy (hunt123):

dont leave yet

OpenStudy (hunt123):

got two more different questions

OpenStudy (hunt123):

A car factory produces 100 cars in a day. If c is the total number of cars produced in d days, which equation represents the number of cars as a function of the number of days?

OpenStudy (unklerhaukus):

i have to sleep, if you close this question, and open a new post someone may help you

OpenStudy (hunt123):

okay thank you so much for your help

OpenStudy (unklerhaukus):

good night!